# ROTATE: Regret-driven Open-ended Training for Ad Hoc Teamwork

### v1

@@ 8 Discussion and Conclusion
This paper reformulates the AHT problem as an open-ended learning problem and introduces ROTATE, a regret-driven algorithm. ROTATE iteratively alternates between improving an AHT agent and generating challenging yet cooperative teammates by optimizing a per-state regret objective designed to discover teammates that exploit cooperative vulnerabilities without encouraging self-sabotage. Empirical evaluations across six AHT tasks demonstrate that ROTATE significantly enhances the generalization capabilities of AHT agents when faced with previously unseen teammates, outperforming a range of baselines from both AHT and UED.
The current work has several limitations that future work may address. First, while this paper provides intuitive justification and strong empirical evidence for the efficacy of the per-state regret objective, it does not provide theoretical analysis. An exciting line of follow-up work is to more formally define the concept of self-sabotage in cooperative MARL and theoretically analyze the properties of the regret objectives proposed here. Second, the paper only studies ROTATE on two-agent, fully observable, and fully cooperative scenarios. Finally, this work has focused on the teammate generation phase of open-ended AHT. Future work might explore ego agent training methods that better handle the nonstationarity induced by open-ended teammate generation.
@@ Appendix A Algorithms
5:$\text{B}_{\pi}\leftarrow\langle\rangle$triangleright Init teammate policy parameter buffer.
@@ A.1 Framework for Open Ended Ad Hoc Teamwork
Section 5 described an open-ended training framework for training an ego agent that can effectively collaborate with previously unseen teammates. We further detail this general open-ended framework in Algorithm 1. In 7, a TeammateGenerator function determines a buffer of teammate policy parameters, $\text{B}^{\text{new}}_{\pi}$. The teammate generator function considers the ego agent’s current policy parameters, $\theta^{\text{ego}}$, and the previous buffer of teammate policy parameters, $\text{B}^{\text{new}}$. Ideally, the teammate generation function generates and samples teammates that induce learning challenges to $\pi^{\text{ego}}$. In 8, an EgoUpdate function specifies a procedure that updates the ego agent’s policy parameters based on the $\text{B}^{\text{new}}_{\pi}$ designed by the teammate generator. Pseudocode for ROTATE, which follows the open-ended framework specified by Algorithm 1, is presented in the following section.
@@ A.2 ROTATE Algorithm
8:$\theta^{-i},\theta^{\text{BR}}\leftarrow\textbf{RandomInit}(\pi),\text{ }\textbf{RandomInit}(\pi)$
9:$\sigma^{\text{BR}}\leftarrow\textbf{RandomInit}(V)$
10:$\sigma^{-i,\text{BR}},\sigma^{-i,\text{ego}}\leftarrow\textbf{RandomInit}(V),\textbf{ RandomInit}(V)$triangleright Init teammate and BR parameters
@@
13:     $s_{\text{XP}}\leftarrow\textbf{SampleStates}(D_{\text{XP}})$triangleright Sample XP states
14:     $D_{\text{SXP}}\leftarrow\textbf{Interact}(\pi_{\theta^{\text{BR}}},\pi_{\theta^{-i}},\mathcal{U}(s_{\text{XP}}))$triangleright Gather SP, XP, and SXP data
15:     $\theta^{\text{BR}}_{\text{old}},\theta^{-i}_{\text{old}}\leftarrow\theta^{\text{BR}},\theta^{-i}$
16:     $\sigma^{\text{BR}}_{\text{old}},\sigma^{-i,\text{BR}}_{\text{old}},\sigma^{-i,\text{ego}}_{\text{old}}\leftarrow\sigma^{\text{BR}},\sigma^{-i,\text{BR}},\sigma^{-i,\text{ego}}$triangleright Store old model parameters.
17:     for $k_{\text{update}}=1,2,\dots,N_{\text{epochs}}$ do
18:         $L_{\text{ppo-clip}}(\theta^{\text{BR}})\leftarrow\textbf{POL\_LOSS\_ADV\_TARG}\left(\theta^{\text{BR}}\theta^{\text{BR}}_{\text{old}},\sigma^{\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}},\epsilon\right)$
19:         $L_{\text{ppo-clip}}(\theta^{-i})\leftarrow\textbf{POL\_LOSS\_ADV\_TARG}\left(\theta^{-i},\theta^{-i}_{\text{old}},\sigma^{-i,\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}},\epsilon\right)$
20:         $L_{\text{reg}}(\theta^{-i})\leftarrow\textbf{POL\_LOSS\_REG\_TARG}\left(\theta^{-i},\theta^{-i}_{\text{old}},\sigma^{-i,\text{BR}}_{\text{old}},\sigma^{-i,\text{ego}}_{\text{old}},D_{\text{XP}},\epsilon\right)$
21:         $L_{V}(\sigma^{\text{BR}})\leftarrow\textbf{VAL\_LOSS}(\sigma^{\text{BR}},\sigma^{\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}})$
22:         $L_{V}(\sigma^{-i,\text{BR}})\leftarrow\textbf{VAL\_LOSS}\left(\sigma^{-i,\text{BR}},\sigma^{-i,\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}}\right)$
23:         $L_{V}(\sigma^{-i,\text{ego}})\leftarrow\textbf{VAL\_LOSS}\left(\sigma^{-i,\text{ego}},\sigma^{-i,\text{ego}}_{\text{old}},D_{\text{XP}}\right)$
24:         $\theta^{\text{BR}}\leftarrow\textbf{GradDesc}(\theta^{\text{BR}},\nabla_{\theta^{\text{BR}}}L_{\text{ppo-clip}}(\theta^{\text{BR}}))$
25:         $\theta^{-i}\leftarrow\textbf{GradDesc}\left(\theta^{-i},\nabla_{\theta^{-i}}\left(L_{\text{ppo-clip}}(\theta^{-i})+L_{\text{reg}}(\theta^{-i})\right)\right)$triangleright Update policies
26:         $\sigma^{\text{BR}}\leftarrow\textbf{GradDesc}(\sigma^{\text{BR}},\nabla_{\sigma^{\text{BR}}}L_{V}(\sigma^{\text{BR}}))$
27:         $\sigma^{-i,\text{BR}}\leftarrow\textbf{GradDesc}(\sigma^{-i,\text{BR}},\nabla_{\sigma^{-i,\text{BR}}}L_{V}(\sigma^{-i,\text{BR}}))$
28:         $\sigma^{-i,\text{ego}}\leftarrow\textbf{GradDesc}(\sigma^{-i,\text{ego}},\nabla_{\sigma^{-i,\text{ego}}}L_{V}(\sigma^{-i,\text{ego}}))$$\triangleright$ Update critics.
29:     end for
@@
31:$\text{B}_{\pi}\leftarrow\text{B}_{\pi}\cup\langle\theta^{-i}\rangle$triangleright Add generated teammate policy parameter
32:Return $\text{B}_{\pi}$
ROTATE’s teammate generation algorithm is detailed in Algorithm 2. As described in Section 6.1, this teammate generation algorithm jointly trains the parameters of a teammate policy and an estimate of its best response (BR) policy, based on a provided ego agent policy. The parameters of the teammate and BR policies, $\theta^{-i}$ and $\theta^{\text{BR}}$, are initialized in 8. The parameters of the BR critic network, $\sigma^{\text{BR}}$, are initialized in 9, while those for the teammate, $\sigma^{-i,\text{BR}}$ and $\sigma^{-i,\text{ego}}$, are initialized in 10. Note that the teammate maintains two critics, for separately estimating returns when interacting with the BR and ego agent policies.
The training of the teammate and BR policies is based on the SP, XP, and SXP interaction data gathered in 12, 13 and 14, which we previously motivated and described in Section 6.1. Recall that an SXP interaction require resetting an environment to start from an available XP state. Since resetting from all available XP states for SXP interaction is impractical, ROTATE samples from XP states to obtain start states for SXP interactions. Experiences from SP, XP, and SXP interaction are stored in buffers $D_{\text{SP}},D_{\text{XP}},D_{\text{SXP}}$ in the form of a collection of tuples, $D=\langle(s_{k},a_{k},r_{k},s^{\prime}_{k})\rangle_{k=1}^{|D|}$. Lines 18 to 28 of Algorithm 2 then highlight how we use the stored experiences to compute loss functions that the trained models optimize.
18 and 19 describe how the teammate and BR policies are trained to mutually maximize returns when interacting with each other during SP and SXP interactions. Both lines call the POL_LOSS_ADV_TARG function, which receives $\left(\theta,\theta_{\text{old}},\sigma_{\text{old}},D,\epsilon\right)$ as input to evaluate the following, standard PPO-clip loss function that encourages return maximization and sufficient exploration:
@@
where $A$ denotes the advantage function. Our implementation of ROTATE uses an estimate of the advantage function obtained via the Generalized Advantage Estimation (GAE) algorithm [47], $A^{\text{GAE}}_{\sigma_{\text{old}}}$. Meanwhile, Line 20 shows how the teammate policy is trained to maximize the ego agent’s regret based on experiences from XP interaction. The POL_LOSS_REG_TARG function that computes a loss function that encourages the maximization of regret is generally the same as the POL_LOSS_ADV_TARG function except for its replacement of the advantage function, $A$, with a regret-based target function defined below:
@@
Rather than optimizing a regret function that requires explicitly computing the return-to-go for SP and XP interaction starting from state $s$, POL_LOSS_REG_TARG estimates the XP return via a 1-step bootstrapped return using the teammate critic parameterized by $\sigma^{-i,\text{BR}}$. Similarly, the SP return is estimated using the teammate critic network parameterized by $\sigma^{-i,\text{ego}}$. This results in a regret optimization method that uses the log-derivative trick to optimize objective functions [58, 19]. The ROTATE regret estimation method and alternative approaches to maximize regret are further discussed in App. C.2.
21, 22 and 23 then detail how we train critic networks that measure returns from the interaction between the generated teammate policy and its best response or ego agent policy. We specifically call the VAL_LOSS function that receives $(\sigma,\sigma_{\text{old}},D)$ to compute the standard mean squared Bellman error (MSBE) loss, defined as:
@@
where $V^{\text{targ}}_{\sigma_{\text{old}}}(s):=A^{\text{GAE}}_{\sigma_{\text{old}}}-V_{\sigma_{\text{old}}}(s)$ is the target value estimate.
The previously defined loss functions can be minimized using any gradient descent-based optimization technique, as we indicate in Lines 24 to 28. In practice, our implementation uses the ADAM optimization technique [26]. At the end of this teammate generation process, 32 and 31 indicate how the generated teammate policy parameter is added to a storage buffer, which is subsequently uniformly sampled to provide teammate policies for ego agent training.
@@
8:$\sigma^{\text{ego}}\leftarrow\textbf{Init}(V)$triangleright Init params of the critic networks of $\pi^{\text{ego}}$
@@
10:     $\theta^{-i}\sim\mathcal{U}(\text{B}_{\pi})$triangleright Sample teammate parameters uniformly
@@
The ego agent policy’s training process proceeds according to Algorithm 3. 10 illustrates how ROTATE creates different teammate policies by uniformly sampling model parameters from the $\text{B}_{\pi}$ resulting from the teammate generation process. Using the experience collaborating with the sampled policies outlined in 11, the ego agent’s policy parameters are updated to maximize its returns via PPO in 14. The only difference between the EGO_POL_LOSS function and POL_LOSS_ADV_TARG function in Algorithm 2 is the input used to compute the loss function. Unlike in the EGO_POL_LOSS function, we assume that the input dataset, $D$, stores the historical sequence of observed states and executed actions, $h$, rather than states. Likewise, we assume that the only difference between the VAL_LOSS and EGO_VAL_LOSS function is that the latter stores the observation-action history rather than states (15). Like recent AHT learning algorithms [64, 40, 38], $\pi^{\text{ego}}$ and $V^{\text{ego}}$ are conditioned on the ego agent’s observation-action history to facilitate an adaptive $\pi^{\text{ego}}$ through an improved characterization of teammates’ policies. The history-conditioned ego architecture and other practical implementation details are described in App. E. Finally, the ego agent update function returns the updated ego agent policy parameters, which are provided as part of the inputs for the next call to ROTATE’s teammate generation function.
@@ Appendix B Baselines Overview
[…]
@@
CoMeDi is a two-stage AHT algorithm. In the first stage, a population of teammates is generated, and in the second stage, an ego agent is trained against the teammate population. The teammate generation stage trains teammate policies one at a time, where the $n$th teammate policy is trained to maximize its SP return, minimize its XP return with the previously generated teammate (i.e. from among teammates $11,\cdots,n-$ 1) that it best collaborates with, and maximizes its “mixed-play" (MP) return. The relationship between the regret objectives described in Section 6 and MP is further discussed in App. C.1.

### v2

@@ 8 Discussion and Conclusion
This paper reformulates AHT as an open-ended learning problem and introduces ROTATE, a regret-driven algorithm. ROTATE iteratively alternates between improving an AHT agent and generating challenging yet cooperative teammates by optimizing a per-state regret objective designed to discover teammates that exploit cooperative vulnerabilities while mitigating self-sabotage. Experiments on an illustrative matrix game demonstrate that the per-state regret objective mitigates self-sabotage. Extensive evaluations across six benchmark tasks demonstrate that ROTATE significantly enhances the generalization capabilities of AHT agents when faced with previously unseen teammates, outperforming baselines from both AHT and UED.
The current work has several limitations. First, while this paper provides intuitive justification and strong empirical evidence for the efficacy of the per-state regret objective, an exciting line of follow-up work is to formally define the concept of self-sabotage and theoretically analyze the properties of the proposed regret objectives. Second, the paper only validates ROTATE on two-agent, fully observable, and fully cooperative scenarios, which leaves the question of whether it scales to more complex scenarios for future work. Finally, this work has focused on the teammate generation phase of open-ended AHT. Future work might explore ego agent training methods that better handle the nonstationarity induced by open-ended teammate generation.
@@ Acknowledgements
Please find the code for this paper at https://github.com/carolinewang01/rotate/.
@@ Appendix A Algorithms
[…]
@@ A.1 Framework for Open Ended Ad Hoc Teamwork
Section 5 described an open-ended training framework for training an ego agent that can effectively collaborate with previously unseen teammates. We further detail this general open-ended framework in Algorithm˜1. In ˜7, a TeammateGenerator function determines a buffer of teammate policy parameters, $\text{B}^{\text{new}}_{\pi}$. The teammate generator function considers the ego agent’s current policy parameters, $\theta^{\text{ego}}$, and the previous buffer of teammate policy parameters, $\text{B}^{\text{new}}$. Ideally, the teammate generation function generates and samples teammates that induce learning challenges to $\pi^{\text{ego}}$. In ˜8, an EgoUpdate function specifies a procedure that updates the ego agent’s policy parameters based on the $\text{B}^{\text{new}}_{\pi}$ designed by the teammate generator. Pseudocode for ROTATE, which follows the open-ended framework specified by Algorithm˜1, is presented in the following section.
@@ A.2 ROTATE Algorithm
[…]
@@
13:  $s_{\text{XP}},s_{\text{SP}}\leftarrow\textbf{SampleStates}(D_{\text{XP}}),\textbf{SampleStates}(D_{\text{SP}})$ $\triangleright$ Sample XP states
14:  $D_{\text{SXP}}\leftarrow\textbf{Interact}(\pi_{\theta^{\text{BR}}},\pi_{\theta^{-i}},\mathcal{U}(s_{\text{XP}}))$
15:  $D_{\text{XSP}}\leftarrow\textbf{Interact}(\pi_{\theta^{\text{ego}}},\pi_{\theta^{-i}},\mathcal{U}(s_{\text{SP}}))$ $\triangleright$ Gather SP, XP, SXP, and XSP data
16:  $\theta^{\text{BR}}_{\text{old}},\theta^{-i}_{\text{old}}\leftarrow\theta^{\text{BR}},\theta^{-i}$
17:  $\sigma^{\text{BR}}_{\text{old}},\sigma^{-i,\text{BR}}_{\text{old}},\sigma^{-i,\text{ego}}_{\text{old}}\leftarrow\sigma^{\text{BR}},\sigma^{-i,\text{BR}},\sigma^{-i,\text{ego}}$ $\triangleright$ Store old model parameters.
18:  for $k_{\text{update}}=1,2,\dots,N_{\text{epochs}}$ do
19:   $L_{\text{ppo-clip}}(\theta^{\text{BR}})\leftarrow\textbf{POL\_LOSS\_ADV\_TARG}\left(\theta^{\text{BR}}\theta^{\text{BR}}_{\text{old}},\sigma^{\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}},\epsilon\right)$
20:   $L_{\text{ppo-clip}}(\theta^{-i})\leftarrow\textbf{POL\_LOSS\_ADV\_TARG}\left(\theta^{-i},\theta^{-i}_{\text{old}},\sigma^{-i,\text{BR}}_{\text{old}},D_{\text{SXP}},\epsilon\right)$
21:   $L_{\text{reg}}(\theta^{-i})\leftarrow\textbf{POL\_LOSS\_REG\_TARG}\left(\theta^{-i},\theta^{-i}_{\text{old}},\sigma^{-i,\text{BR}}_{\text{old}},\sigma^{-i,\text{ego}}_{\text{old}},D_{\text{SP}}\cup D_{\text{XP}},\epsilon\right)$
22:   $L_{V}(\sigma^{\text{BR}})\leftarrow\textbf{VAL\_LOSS}(\sigma^{\text{BR}},\sigma^{\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}})$
23:   $L_{V}(\sigma^{-i,\text{BR}})\leftarrow\textbf{VAL\_LOSS}\left(\sigma^{-i,\text{BR}},\sigma^{-i,\text{BR}}_{\text{old}},D_{\text{SP}}\cup D_{\text{SXP}}\right)$
24:   $L_{V}(\sigma^{-i,\text{ego}})\leftarrow\textbf{VAL\_LOSS}\left(\sigma^{-i,\text{ego}},\sigma^{-i,\text{ego}}_{\text{old}},D_{\text{XP}}\cup D_{\text{XSP}}\right)$
25:   $\theta^{\text{BR}}\leftarrow\textbf{GradDesc}(\theta^{\text{BR}},\nabla_{\theta^{\text{BR}}}L_{\text{ppo-clip}}(\theta^{\text{BR}}))$
26:   $\theta^{-i}\leftarrow\textbf{GradDesc}\left(\theta^{-i},\nabla_{\theta^{-i}}\left(L_{\text{ppo-clip}}(\theta^{-i})+L_{\text{reg}}(\theta^{-i})\right)\right)$ $\triangleright$ Update policies
27:   $\sigma^{\text{BR}}\leftarrow\textbf{GradDesc}(\sigma^{\text{BR}},\nabla_{\sigma^{\text{BR}}}L_{V}(\sigma^{\text{BR}}))$
28:   $\sigma^{-i,\text{BR}}\leftarrow\textbf{GradDesc}(\sigma^{-i,\text{BR}},\nabla_{\sigma^{-i,\text{BR}}}L_{V}(\sigma^{-i,\text{BR}}))$
29:   $\sigma^{-i,\text{ego}}\leftarrow\textbf{GradDesc}(\sigma^{-i,\text{ego}},\nabla_{\sigma^{-i,\text{ego}}}L_{V}(\sigma^{-i,\text{ego}}))\triangleright$ Update critics.
@@
31:end for
[…]
33:Return $\text{B}_{\pi}$
ROTATE’s teammate generation algorithm is detailed in Algorithm 2. As described in Section 6.1, this teammate generation algorithm jointly trains the parameters of a teammate policy and an estimate of its best response (BR) policy, based on a provided ego agent policy. The parameters of the teammate and BR policies, $\theta^{-i}$ and $\theta^{\text{BR}}$, are initialized in ˜8. The parameters of the BR critic network, $\sigma^{\text{BR}}$, are initialized in ˜9, while those for the teammate, $\sigma^{-i,\text{BR}}$ and $\sigma^{-i,\text{ego}}$, are initialized in ˜10. Note that the teammate maintains two critics, for separately estimating returns when interacting with the BR and ego agent policies.
The training of the teammate and BR policies is based on the SP, XP, XSP, and SXP interaction data gathered in ˜12, 13, 14 and 15, which we previously motivated and described in Section 6.1. Recall that an SXP interaction require resetting an environment to start from an available XP state, and an XSP interaction analogously requires resetting to an SP state. Since resetting from all available XP states for SXP interaction is impractical, ROTATE samples from XP states to obtain start states for SXP interactions (and similarly for XSP). Experiences from SP, XP, SXP, and XSP interaction are stored in buffers $D_{\text{SP}},D_{\text{XP}},D_{\text{SXP}},D_{\text{XSP}}$ in the form of a collection of tuples, $D=\langle(s_{k},a_{k},r_{k},s^{\prime}_{k})\rangle_{k=1}^{|D|}$. Lines 19 to 29 of Algorithm 2 then highlight how we use the stored experiences to compute loss functions that the trained models optimize.
˜19 and 20 describe how the teammate and BR policies are trained to mutually maximize returns when interacting with each other during SP and SXP interactions. Both lines call the POL_LOSS_ADV_TARG function, which receives $\left(\theta,\theta_{\text{old}},\sigma_{\text{old}},D,\epsilon\right)$ as input to evaluate the following, standard PPO-clip loss function that encourages return maximization and sufficient exploration:
@@
where $A$ denotes the advantage function. Our implementation of ROTATE uses an estimate of the advantage function obtained via the Generalized Advantage Estimation (GAE) algorithm [47], $A^{\text{GAE}}_{\sigma_{\text{old}}}$. Meanwhile, Line 21 shows how the teammate policy is trained to maximize the ego agent’s regret based on experiences from XP and SP interactions. The POL_LOSS_REG_TARG function that computes a loss function that encourages the maximization of regret is generally the same as the POL_LOSS_ADV_TARG function except for its replacement of the advantage function, $A$, with a regret-based target function. The regret-based target function is defined differently but symmetrically for SP and XP states. We describe the target function for regret from XP states below, and refer the reader to the code for the target function for regret from SP states.
@@
Rather than optimizing a regret function that requires explicitly computing the return-to-go, POL_LOSS_REG_TARG estimates the XP return via a 1-step bootstrapped return using the teammate critic parameterized by $\sigma^{-i,\text{ego}}$. Similarly, the SP return is estimated using the teammate critic network parameterized by $\sigma^{-i,\text{BR}}$. This results in a regret optimization method that uses the log-derivative trick to optimize objective functions [58, 21]. The ROTATE regret estimation method and alternative approaches to maximize regret are further discussed in App. C.2.
˜22, 23 and 24 then detail how we train critic networks that measure returns from the interaction between the generated teammate policy and its best response or ego agent policy. We specifically call the VAL_LOSS function that receives $(\sigma,\sigma_{\text{old}},D)$ to compute the standard mean squared Bellman error (MSBE) loss, defined as:
@@
where $V^{\text{targ}}_{\sigma_{\text{old}}}(s):=A^{\text{GAE}}_{\sigma_{\text{old}}}+V_{\sigma_{\text{old}}}(s)$ is the target value estimate.
The previously defined loss functions can be minimized using any gradient descent-based optimization technique, as we indicate in Lines 25 to 29. In practice, our implementation uses the ADAM optimization technique [26]. At the end of this teammate generation process, ˜33 and 32 indicate how the generated teammate policy parameter is added to a storage buffer, which is subsequently uniformly sampled to provide teammate policies for ego agent training.
@@
[…]
@@
10:  $\theta^{-i}\sim\mathcal{U}(\text{B}_{\pi})$ $\triangleright$ Sample teammate parameters uniformly
@@
The ego agent policy’s training process proceeds according to Algorithm 3. ˜10 illustrates how ROTATE creates different teammate policies by uniformly sampling model parameters from the $\text{B}_{\pi}$ resulting from the teammate generation process. Using the experience collaborating with the sampled policies outlined in ˜11, the ego agent’s policy parameters are updated to maximize its returns via PPO in ˜14. The only difference between the EGO_POL_LOSS function and POL_LOSS_ADV_TARG function in Algorithm 2 is the input used to compute the loss function. Unlike in the EGO_POL_LOSS function, we assume that the input dataset, $D$, stores the historical sequence of observed states and executed actions, $h$, rather than states. Likewise, we assume that the only difference between the VAL_LOSS and EGO_VAL_LOSS function is that the latter stores the observation-action history rather than states (˜15). Like recent AHT learning algorithms [64, 40, 39], $\pi^{\text{ego}}$ and $V^{\text{ego}}$ are conditioned on the ego agent’s observation-action history to facilitate an adaptive $\pi^{\text{ego}}$ through an improved characterization of teammates’ policies. The history-conditioned ego architecture and other practical implementation details are described in App. E. Finally, the ego agent update function returns the updated ego agent policy parameters, which are provided as part of the inputs for the next call to ROTATE’s teammate generation function.
@@ Appendix B Baselines Overview
A two-stage AHT algorithm where a population of “confederate" and best-response agent pairs is generated, and an ego agent is trained against the confederates. BRDiv maintains a cross-play matrix containing the returns for all confederate and best-response pairs. The diagonal returns (self-play) are maximized, while the off-diagonal returns (cross-play) are minimized. BRDiv and LIPO [10] share a similar objective, where the main differences are: (1) If xp_weight denotes the weight on the XP return, then BRDiv requires that the coefficient on the SP return is always $1+2*\texttt{xp\_weight}$, and (2) LIPO introduces a secondary diversity metric based on mutual information, and (3) LIPO assumes that agents within a team (i.e., a confederate-BR pair) share parameters.
@@
CoMeDi is a two-stage AHT algorithm. In the first stage, a population of teammates is generated, and in the second stage, an ego agent is trained against the teammate population. The teammate generation stage trains teammate policies one at a time, where the $n$th teammate policy is trained to maximize its SP return, minimize its XP return with the previously generated teammate (i.e. from among teammates $1,\cdots,n-1$) that it best collaborates with, and maximizes its “mixed-play" (MP) return. The relationship between the regret objectives described in Section 6 and MP is further discussed in App. C.1.